OpenStudy (anonymous):

simplify each expression: 4x^2y^5 -------- 12x^3y^2z^-3

OpenStudy (anonymous):

\[\frac{4 x^2 y^5}{\left(\frac{12 x^3 y^2}{z^3}\right)}\text{=}\frac{y^3 z^3}{3 x}\]